OpenStudy (anonymous):

raymond is buying boxes of party favors for his son's birthday he has to decide how many boxes he should purchase. twenty eight children will attend the birthday party. if each child gets 10 party favors and there are 70 favors in each box, how many boxes should raymond buy ? please help !

OpenStudy (anonymous):

28*10=280 boxes so multiply 70 until you get 280

OpenStudy (anonymous):

Like this 70*1=70 70*2=140 70*3=210 70*4=

OpenStudy (anonymous):

ohh yeah that makes sense, thank you so much ! :)

OpenStudy (anonymous):

Your welcome :)